OpenStudy (mar1977):

simplify 6^2 over 6^4 36 1/12 1/36 66 I'm think 1/36 but it seems I'm doing something wrong any help

OpenStudy (steve816):

Know your exponent rules buddy!

OpenStudy (tkhunny):

Show us what you did. We will then tell you that it is correct if that was the result.

OpenStudy (anonymous):

You are thinking correct 6*6=36 6*6*6*6=1296 36/1296=0.02777777777 1/36=0.02777777777

OpenStudy (anonymous):

@mar1977

OpenStudy (tkhunny):

{sigh} Now, we'll never know.

OpenStudy (mathstudent55):